OpenStudy (anonymous):

If the point (3/5,4/5) corresponds to an angle 0 in the unit circle, what is csc 0 ?

OpenStudy (anonymous):

sin theta = 4/5 so, csc theta = 1/sin theta = 5/4

OpenStudy (anonymous):

so is it 4/5 or 5/4 because I have both those options for my multiple choice

OpenStudy (anonymous):

5/4 is final answer. I just ave you a reason for my answer.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

If you are riding on a Ferris wheel that is 60 feet tall and it takes you 5 minutes to make one revolution, what is your linear velocity?